The displacement of a harmonic oscillator is given by x(t)=9.4 sin (15t), where the unit of x are

meters and t is measures in seconds.

(a) Find the amplitude and the frequency. (b) What is the maximum velocity of the motion?

(c) What is the maximum acceleration?


1
Expert's answer
2021-01-14T10:38:20-0500

(a) The amplitude is "A = 9.4\ m".

We can find the frequency from the formula:


"\omega=2\pi f,""f=\dfrac{\omega}{2\pi}=\dfrac{15\ \dfrac{rad}{s}}{2\pi}=2.38\ Hz."

(b) Let’s take the derivative of "x(t)" with respect to time:


"v(t)=\dfrac{d}{dt}(9.4sin(15t))=9.4\ m\cdot15\ \dfrac{rad}{s}\cdot cos(15t),""v(t)=141\ \dfrac{m}{s}\cdot cos(15t)."

We can obtain the maximum velocity of the motion when "cos(\omega t)=1", therefore:


"v(t)_{max}=141\ \dfrac{m}{s}."


(c) Let’s take the derivative of "v(t)" with respect to time:


"a(t)=\dfrac{d}{dt}(141\ \dfrac{m}{s}\cdot cos(15t))=-141\ \dfrac{m}{s}\cdot15\ \dfrac{rad}{s}sin(15t),""a(t)=-2115\ \dfrac{m}{s^2}\cdot sin(15t)."

We can obtain the maximum acceleration of the motion when "sin(\omega t)=1", therefore:


"a(t)_{max}=-2115\ \dfrac{m}{s^2}."

Answer:

(a) "A = 9.4\ m, f=2.38\ Hz."

(b) "v(t)_{max}=141\ \dfrac{m}{s}."

(c) "a(t)_{max}=-2115\ \dfrac{m}{s^2}."